OpenStudy (anonymous):

Karim borrowed one hundred twenty dollars from his older brother for the purchase of a skateboard. He promised to pay back the loan in 6 months at a simple annual interest rate of fifteen percent. How much will Karim owe his brother at that time?

OpenStudy (anonymous):

@djcool31

OpenStudy (***[isuru]***):

annual interest for 120 dollars = (15/ 100) x 120 then the monthly interest = (15/ 100) x 120 x (1/12) [ dividing the annual interest by 12 ] total interest to pay = (15/ 100) x 120 x (1/12) x 6 [ multiply by 6 months ] total amount to pay = total interest to pay + 120 dollars can u simplify it from here ?

OpenStudy (anonymous):

@martin103